OpenStudy (anonymous):

Integrate: 1+sinx/cos^2x

OpenStudy (raden):

first, change the equation by using some identities of trigono : (1+sinx)/cos^2x = 1/cos^2x + sinx/cos^2x = sec^2x + sinx/cosx * 1/cosx = sec^2x + tanx sec x see the integral becomes